OpenStudy (anonymous):

how to solve f(x,y) = 4x^3 + y^2 ; 2x^2 + y^2 = 1 using lagrange multipliers????

OpenStudy (anonymous):

delf = λdelg where g = 2x^2 + y^2 -1 (12x^2,2y) =λ(4x,2y) 2y = 2λy 2y(λ-1) = 0 λ = 1,y=0 λ = 1: 12x^2 = 4xλ 4x(λ-3x)=0 4x(1-3x)=0 x = 0, x = 1/3 so for x = 0 plug into 2x^2 + y^2 = 1 y= +/-1 for x = 1/3 plug into 2x^2 + y^2 = 1 y = +/- sqrt(7/9) y = 0 x = +/- sqrt(2)

OpenStudy (anonymous):

thank you so much!!!! ;0)

OpenStudy (anonymous):

How would I solve : find the point on the line 2x-4y = 3 that is closest to hte origin using the same method?